A buffer solution is prepared by dissolving 1.50 g each of benzoic acid, C7H6O2, and

sodium benzoate, NaC7H5O2, in 150.0 mL of solution. [Ka = 6.3 x 10-5]

a. What is the pH of this buffer solution?

b. Which buffer component must be added, and in what quantity, to change the

buffer pH to 4.00?

c. What quantity of 2.0 M NaOH or 2.0 M HCl must be added to the original buffer to

change the pH to 4.00.


1
Expert's answer
2022-03-03T08:34:02-0500

Henderson-Hasselbalch Equation 𝑝𝐻 = π‘πΎπ‘Ž + π‘™π‘œπ‘”"\frac{[A-]}{[HA]}"

𝑝𝐻 = βˆ’ log(6.3 "\times" 10 βˆ’5 ) + π‘™π‘œπ‘” "\frac{(0.015)} {(0.150)}"

pH = 4.1 + 0.10 = 4.20